a proton undergo a head on elastic collision with a particle of unknown mass which is initially at rest and rebounds with 4/9 of its initial kinetic energy. Calculate the ratio of the unknown mass with respect to the mass of proton.
1
Expert's answer
2019-10-18T10:16:17-0400

From the conservation of momentum:


"mv=-mv'+Mu"

We have:

"0.5mv'^2=\frac{4}{9}(0.5mv^2)\to v'=\frac{2}{3}v"

So,


"mv=-m\frac{2}{3}v+Mu\to u=v\frac{5}{3}\frac{m}{M}"

From the conservation of energy:


"0.5Mu^2=0.5mv^2-0.5mv'^2=\frac{5}{9}(0.5mv^2)"

"M\left(v\frac{5}{3}\frac{m}{M}\right)^2=\frac{5}{9}(mv^2)"


"\left(\frac{5}{3}\right)^2=\frac{5}{9}\frac{M}{m}"

"\frac{M}{m}=5"